# Analysis of a finite element method for second order uniformly elliptic PDEs in non-divergence form

### v1

@@ 3.1. Global $W^{2,p}$ estimate of linear elliptic PDE in non-divergence form with uniformly continuous AA
$-\nabla\cdot\left(A_{0}\nabla M(f)\right)=f,\qquad\forall f\in L^{\tilde{q}}(\Omega),\forall 1<\tilde{q}\leq q.$
@@
$Tw=\left(M\tilde{w}\right)|_{B_{2R}(\boldsymbol{x}_{0})}$
@@
$\tilde{w}=(A_{0}-A):D^{2}w\text{ in }B_{2R}(\boldsymbol{x}_{0}),\qquad\tilde{w}=0\text{ in }\Omega\setminus\overline{B_{2R}(\boldsymbol{x}_{0})}.$
@@
$w=Tw+M\left(A:D^{2}v\right)|_{B_{2R_{0}(\boldsymbol{x}_{0})}}\text{ in }B_{2R_{0}(\boldsymbol{x}_{0})}.$
On the other hand, since $v=0$ in $\Omega\setminus\overline{B_{2R}(\boldsymbol{x}_{0})}$ and $v\in W^{2,p}(\Omega)$ ($1<p<q$), we have that
$v|_{B_{2R_{0}}(\boldsymbol{x}_{0})}=T(v|_{B_{2R_{0}}(\boldsymbol{x}_{0})})+M\left(A:D^{2}v\right)|_{B_{2R_{0}}(\boldsymbol{x}_{0})}\text{ in }B_{2R_{0}}(\boldsymbol{x}_{0}).$
Therefore we have that $\phi|_{B_{R_{0}}(\boldsymbol{x}_{0})}\in W^{2,q}(B_{R_{0}}(\boldsymbol{x}_{0}))$.
@@ 3.2. Stability analysis of FEM for linear elliptic PDE in non-divergence form with uniformly continuous AA
(28, 29) imply that for any $p\in(1,+\infty]$,
@@
Then by (33,34) and the discrete inverse inequality, we can conclude that the proof is complete. ∎
@@ 3.2.2. Stability estimates for linear elliptic PDE in non-divergence form with uniformly continuous AA
Therefore, (42,44) imply that
@@
We notice $I_{h}(\eta v_{h})\in V_{h}(B_{\frac{11}{6}R})$. Then by (36,8,40,43), we have
@@
for a subsequence of $h\rightarrow 0$. (52, 54) imply that $\nabla w=\boldsymbol{\phi}$ almost everywhere in $\Omega$. Therefore, we have that $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ and
@@
According to (55, 56, 58), We can conclude the proof is complete. ∎
@@
We divide the proof into two steps. In step $1$, We will first prove the estimates (67) and (68,69) for $1<p\leq 2$ if $\Omega$ is convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), and $\frac{4}{3}-\epsilon_{1}<p<\frac{4}{3}+\epsilon_{2}$ if $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$. In step $2$, we will prove the estimate (67) for $1<p\leq\frac{4}{3}-\epsilon_{1}$ and $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$.
@@ 4. Hamilton-Jacobi-Bellman equation
For the elliptic HJB equation (3), the main theoretical results are Theorem 4.1 and Theorem 4.3. Theorem 4.1 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ of the elliptic HJB equation (3), while Theorem 4.3 provides optimal convergence of the FEM (16).
@@
In addition, if (4,5) hold and $B\in[L^{p}(\Omega)]^{d\times d},\boldsymbol{c}\in[L^{p}(\Omega)]^{d},w\in L^{p}(\Omega)$ and $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$ for some $p\in[1,+\infty)$, then we have that $\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]\in L^{p}(\Omega)$ and a positive constant $C$ satisfying
@@
According to (74), (4,5) and the fact $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$ imply that $\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]\in L^{p}(\Omega)$ and (75) holds. ∎
@@
On the other hand, (76a,77a, 77b) implies
@@ 4.1. Well-posedness of Hamilton-Jacobi-Bellman equation
Let $\Omega$ be a Lipschitz polyhedral in $\mathbb{R}^{d}$ ($d=2,3$). We assume that $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$, and (4,5) and the assumption (6) hold. There is a constant $0\leq\underline{\kappa}<1$ which may depend on $p$, such that if
@@
By (4,5) and the fact $\tilde{\Lambda}$ is a countable set, there is a set $E_{1}\supset E_{0}$ with zero $d$-dimensional Lebesgue measure such that
@@
Obviously, the assumption (6) is satisfied by $(\gamma^{\lambda,\alpha}A^{\alpha},\gamma^{\lambda,\alpha}\boldsymbol{b}^{\alpha},\gamma^{\lambda,\alpha}c^{\alpha},\gamma^{\lambda,\alpha}f^{\alpha})$ with $\tilde{\Lambda}\subset\Lambda$ and the set $E_{1}$. Then by Lemma 4.1 and (4,5), we have that for any $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$,
@@ 4.2. Optimal convergence of numerical solution of FEM for Hamilton-Jacobi-Bellman equation
Let $\Omega$ be a Lipschitz polyhedral in $\mathbb{R}^{d}$ ($d=2,3$). We assume that $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$, and (4,5) and the assumption (6) hold. There is a constant $0\leq\kappa<1$ which my depend on $p$, such that if
@@
Since (4,5) and the assumption (6) hold, it is obvious to mimic the proof of Theorem 4.1 (the argument to obtain (83)) to have that the finite element method (16) is equivalent to find $u_{h}\in V_{h}$ satisfying

### v2

@@ 3.1. Global $W^{2,p}$ estimate of linear elliptic PDE in non-divergence form with uniformly continuous AA
$\nabla\cdot\left(A_{0}\nabla M(f)\right)=f,\qquad\forall f\in L^{\tilde{q}}(\Omega),\forall 1<\tilde{q}\leq q.$
@@
(22) $Tw=\left(M\tilde{w}\right)|_{B_{2R}(\boldsymbol{x}_{0})}$
@@
(23) $\tilde{w}=(A_{0}-A):D^{2}w\text{ in }B_{2R}(\boldsymbol{x}_{0}),\qquad\tilde{w}=0\text{ in }\Omega\setminus\overline{B_{2R}(\boldsymbol{x}_{0})}.$
@@
(24) $w=Tw+M\left(A:D^{2}v\right)|_{B_{2R_{0}(\boldsymbol{x}_{0})}}\quad\text{ in }B_{2R_{0}(\boldsymbol{x}_{0})}.$
On the other hand, since $v=0$ in $\Omega\setminus\overline{B_{2R}(\boldsymbol{x}_{0})}$ and $v\in W^{2,p}(\Omega)$ ($1<p<q$), (22) and (23) imply that
$T(v|_{B_{2R_{0}}(\boldsymbol{x}_{0})})=M\left((A_{0}-A):D^{2}v\right)|_{B_{2R_{0}}(\boldsymbol{x}_{0})}.$
Then we have
(25) $v|_{B_{2R_{0}}(\boldsymbol{x}_{0})}=M(A_{0}:D^{2}v)|_{B_{2R_{0}}(\boldsymbol{x}_{0})}$
$=$ $T(v|_{B_{2R_{0}}(\boldsymbol{x}_{0})})+M\left(A:D^{2}v\right)|_{B_{2R_{0}}(\boldsymbol{x}_{0})}\quad\text{ in }B_{2R_{0}}(\boldsymbol{x}_{0}).$
(24) and (25) imply that $\phi|_{B_{R_{0}}(\boldsymbol{x}_{0})}\in W^{2,q}(B_{R_{0}}(\boldsymbol{x}_{0}))$.
@@ 3.2. Stability analysis of FEM for linear elliptic PDE in non-divergence form with uniformly continuous AA
(32, 33) imply that for any $p\in(1,+\infty]$,
@@
Then by (37,38) and the discrete inverse inequality, we can conclude that the proof is complete. ∎
@@ 3.2.2. Stability estimates for linear elliptic PDE in non-divergence form with uniformly continuous AA
Therefore, (46,48) imply that
@@
We notice $I_{h}(\eta v_{h})\in V_{h}(B_{\frac{11}{6}R})$. Then by (40,8,44,47), we have
@@
for a subsequence of $h\rightarrow 0$. (56, 58) imply that $\nabla w=\boldsymbol{\phi}$ almost everywhere in $\Omega$. Therefore, we have that $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ and
@@
According to (59, 60, 62), We can conclude the proof is complete. ∎
@@
We divide the proof into two steps. In step $1$, We will first prove the estimates (71) and (72,73) for $1<p\leq 2$ if $\Omega$ is convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), and $\frac{4}{3}-\epsilon_{1}<p<\frac{4}{3}+\epsilon_{2}$ if $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$. In step $2$, we will prove the estimate (71) for $1<p\leq\frac{4}{3}-\epsilon_{1}$ and $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$.
@@ 4. Hamilton-Jacobi-Bellman equation
For the elliptic HJB equation (2), the main theoretical results are Theorem 4.1 and Theorem 4.3. Theorem 4.1 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ of the elliptic HJB equation (2) if the coefficients satisfy the condition (83), while Theorem 4.3 provides optimal convergence of the FEM (16) if the coefficients satisfy the condition (93).
4.1. Auxiliary results for HJB equation
@@
In addition, if (3,4) hold and $B\in[L^{p}(\Omega)]^{d\times d},\boldsymbol{c}\in[L^{p}(\Omega)]^{d},w\in L^{p}(\Omega)$ and $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$ for some $p\in[1,+\infty)$, then we have that $\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]\in L^{p}(\Omega)$ and a positive constant $C$ satisfying
@@
According to (78), (3,4) and the fact $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$ imply that $\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]\in L^{p}(\Omega)$ and (79) holds. ∎
@@
On the other hand, (80a,81a, 81b) implies
@@ 4.2. Well-posedness of Hamilton-Jacobi-Bellman equation
When the domain is convex and $p=2$, the condition (83) for coefficients in Theorem 4.1 is identical to the Cordès condition used in [40].
@@
Let $\Omega$ be a Lipschitz polyhedral in $\mathbb{R}^{d}$ ($d=2,3$). We assume that $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$, and (3,4) and the assumption (6) hold. There is a constant $0\leq\underline{\kappa}<1$ which may depend on $p$, such that if
@@
By (3,4) and the fact $\tilde{\Lambda}$ is a countable set, there is a set $E_{1}\supset E_{0}$ with zero $d$-dimensional Lebesgue measure such that
@@
Obviously, the assumption (6) is satisfied by $(\gamma^{\lambda,\alpha}A^{\alpha},\gamma^{\lambda,\alpha}\boldsymbol{b}^{\alpha},\gamma^{\lambda,\alpha}c^{\alpha},\gamma^{\lambda,\alpha}f^{\alpha})$ with $\tilde{\Lambda}\subset\Lambda$ and the set $E_{1}$. Then by Lemma 4.1 and (3,4), we have that for any $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$,
@@ 4.3. Optimal convergence of numerical solution of FEM for Hamilton-Jacobi-Bellman equation
When the domain is convex and $p=2$, the condition (93) for coefficients in Theorem 4.3 is more restrictive than the Cordès condition used in [40]. The stronger restriction of the condition (93) is due to two reasons. The first reason is that the constant of the discrete Miranda–Talenti inequality of the discrete Laplacian operator $L_{I,h}$ used in the analysis of the FEM (16) is not known explicitly and probably bigger than $1$. The second reason is that we don’t want to include the parameter $\lambda$ used in [40], which may not be known beforehand, in the design of the weight function $\gamma^{\alpha}$ (see the difference between (15) and the weight function used in [40]). Though it is more restrictive, Remark 4.5 shows that the condition (93) allows non-trivial examples. On the other hand, Theorem 4.3 is valid for a wide range of $p$ (see (5b)) and is valid on two dimensional non-convex polygon for a neighbourhood of $\frac{4}{3}$.
@@
Let $\Omega$ be a Lipschitz polyhedral in $\mathbb{R}^{d}$ ($d=2,3$). We assume that $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$, and (3,4) and the assumption (6) hold. There is a constant $0\leq\kappa<1$ which my depend on $p$, such that if
@@
Since (3,4) and the assumption (6) hold, it is obvious to mimic the proof of Theorem 4.1 (the argument to obtain (87)) to have that the finite element method (16) is equivalent to find $u_{h}\in V_{h}$ satisfying